# Spiffy: Multiplying Diffusion LLM Acceleration via Lossless Speculative Decoding

### v2

@@ 4.2 Verification
Since the overheads of Spiffy are negligible compared to the model inference time (Appendix B), this value will be close to the true speedup achieved by the algorithm. Moreover, we show that Algorithm 1 is lossless, that is, it preserves the output distribution of the model. We include a proof of this statement in Appendix A.1 as well as numerical experiments in Appendix A.2 to demonstrate this property.
@@ 5.1 Main Experiment
Additionally, we see that model accuracy on downstream metrics is preserved up to minor differences in generations caused due to precision errors. These precision discrepancies are attributed to torch.bfloat16 GPU matrix multiplication accumulation errors for tensors of different sizes and are discussed further in Appendix A.2.
@@
Table 1: Spiffy achieves speedups of up to $3\times$ for various dLLMs and tasks, while preserving accuracy on downstream metrics. Speedup :=:= (Originally Required NFEs)/(Final NFEs). Exact generation is preserved up to numerical errors caused by BF16 matrix multiplication.
@@ A.2 Experimental Verification
We further validate our claim above by computing the downstream accuracy of the chosen models with and without Spiffy and comparing their resulting metric scores. These results are displayed in 1. Each task defined by lm-eval has an associated metric which is computed automatically along with an associated standard error. Table 1 demonstrates that Spiffy is indeed a lossless method with respect to the vanilla model inference.
Minor differences in the metric values, of the order $0.01$, are attributed to minor variations in generated outputs due to precision loss caused by torch.bfloat16 matrix multiplication on CUDA GPUs. This is caused due to differing accumulation orders for tensors of differing sizes. In particular, even though each of the $10$ draft blocks in Spiffy with $D=10$ are independent, matrix multiplications with this extended tensor will have a slightly different accumulation order as compared to an input tensor with only $D=3$ draft blocks added to it. This leads to minor differences in generations and due to cumulative nature of errors, may lead to different final results for tasks that compute exact_match or pass@1 as their metric. For applications where precision is critical, we recommend padding the input tensor to $4096$ to ensure predictable accumulation on CUDA GPU’s in reduced precision. We do not adopt this method here as it is prohibitively expensive and minor differences do not affect the output significantly as demonstrated by Table 1. The PyTorch documentation on numerical accuracy may be referred to for further explanation on floating point error accumulation in such settings.

### v3

@@ 4.2 Verification
Since the overheads of Spiffy are negligible compared to the model inference time (Appendix B), this value will be close to the true speedup achieved by the algorithm. Moreover, we show that Algorithm 1 is lossless, that is, it preserves the output distribution of the model. We include a proof of this statement in Appendix A.1.
While Algorithm 1 is lossless in general, in certain settings, producing the exact draft probabilities $\{p^{m}_{\text{DD}}\}$ is computationally expensive. In particular, for full-sequence bidirectional decoding, achieving exact decoding requires construction of a batched attention mask to ensure that all the keys and values attended to by the draft blocks, in turn, attend to the draft blocks. This would increase the computational requirements of the system by a factor proportional to the number of draft blocks and is hence infeasible. Instead, we choose to create a simple blockwise tree-attention mask structure that we see achieves near-lossless performance for full-sequence bidirectional dLLMs such as LLaDA in its default inference setting. Moreover, using this simple mask, we can achieve exact lossless speculative decoding when utilizing dual caches (Wu et al. 2025) or when decoding blockwise causal dLLMs such as SDAR (Cheng et al. 2025). In this work, we demonstrate near-lossless decoding on LLaDA models, which provides an upper bound on the error of our method, through numerical experiments in Appendix A.2.
@@ 5.1 Main Experiment
Additionally, we see that model accuracy on downstream metrics is preserved up to minor differences in generations caused due to fully-bidirectional decoding of LLaDA as noted in Sec 4.2 and precision errors. These precision discrepancies are attributed to torch.bfloat16 GPU matrix multiplication accumulation errors for tensors of different sizes and are discussed further in Appendix A.2.
@@
Table 1: Spiffy achieves speedups of up to $3\times$ for various dLLMs and tasks, while preserving accuracy on downstream metrics. Speedup :=:= (Originally Required NFEs)/(Final NFEs).
@@ A.2 Experimental Verification
We note that achieving the losslessness described above requires generating draft probabilities such that $p_{\text{DD}}^{m}(t_{0}-2)=p_{\text{TD}}(t_{0}-2)$. That is, the output distribution of the draft block should be identical to the distribution produced by the same block during true vanilla inference. In some settings, ensuring this is computationally expensive because the draft block must attend to the same information as the true block, but must also attend to the original block in order to generate the target distribution. Concretely, suppose we have 3 blocks $X_{1},X_{2},X_{3}$ and we are currently decoding $X_{2}(t)$ with a draft block $X_{2}^{*}(t-1)$. Since $X_{2}^{*}$ attends to the entire sequence (excluding $X_{2}$), but $X_{1}$ and $X_{3}$ in turn attend to $X_{2}$, the draft block is indirectly attending to the current block at a previous timestep $t$. This leakage of information would lead to drift in the output hidden states of the draft block across layers. However, two factors mitigate this deviation. First, $X_{1}$ has already been fully decoded in semi-autoregressive inference, so its hidden states remain stable during decoding (Wu et al. 2025). Second, since $X_{3}$ consists entirely of masked tokens, both $X_{2}$ and $X_{2}^{*}$ place a low attention mass on this block. Due to these factors, the shift in output distribution is minimal.
We validate our claim above by computing the downstream accuracy of the LLaDA model with full-sequence bidirectional decoding with and without Spiffy and comparing their resulting metric scores. These results are displayed in Table 1. Each task defined by lm-eval has an associated metric which is computed automatically along with an associated standard error. Table 1 demonstrates that even in the case of fully-bidirectional decoding, Spiffy is indeed a lossless method with respect to the vanilla model inference. In this framework, we may also conclude that when a dLLM is run with blockwise causal masking (Cheng et al. 2025), the previous hidden states do not attend to future hidden states, which will ensure that the draft and true distribution remain identical. Similarly, if a dLLM is run with dual KV caches, before and after the current block (Wu et al. 2025), the keys and values of $X_{1}$ and $X_{3}$ remain constant, resulting in an exact draft distribution.
We further investigate minor differences in the metric values, on the order $0.01$, and attribute them to minor variations in generated outputs due to precision loss caused by torch.bfloat16 matrix multiplication on CUDA GPUs. This is caused due to differing accumulation orders for tensors of differing sizes. In particular, even though each of the $10$ draft blocks in Spiffy with $D=10$ are independent, matrix multiplications with this extended tensor will have a slightly different accumulation order as compared to an input tensor with only $D=3$ draft blocks added to it. This leads to minor differences in generations and due to cumulative nature of errors, may lead to different final results for tasks that compute exact_match or pass@1 as their metric. For applications where precision is critical, we recommend padding the input tensor to $4096$ to ensure predictable accumulation on CUDA GPU’s in reduced precision. We do not adopt this method here as it is prohibitively expensive and minor differences do not affect the output significantly as demonstrated by Table 1. The PyTorch documentation on numerical accuracy may be referred to for further explanation of floating point error accumulation in such settings.